tos .J 7 Formulate the Newtons method for +2x y+2z J&z=-F).
Answer #1

y-11 f(x, y)-

Then 2.T f(x, y) = y + 2 + 1 + 2x-1, g(x, y) =

That is, f(x,y)=\frac{y}{x}+\frac{2x}{y}+2,\,g(x,y)=\frac{x+y}{y+2x}-11

And so 12(r y) (y+2x)2, gy(x,y)= 1 2r y+2x

So the Jacobian is

12a r+y 2(r+y1

And so the required iterative system is

X_{n+1}=X_n-\left(J(X_n) \right )^{-1}F(X_n) which is

1 12 with (X_0,Y_0)=(1,1)